### v1

@@ 5.1 Design of Curriculum Learning in Diffusion Models
As we observed in Section 4, difficulties in denoising tasks increase as $t$ gets smaller. To utilize this in our training strategy for diffusion models, we adopt an easy-to-hard curriculum learning approach by partitioning the entire range of timesteps into $N$ clusters, denoted as $\{C_{i}\}_{i=1}^{N}$, where each cluster $C_{i}$ spans an interval $[l_{i},r_{i}]$, ensuring $l_{i}<r_{i}$ and continuity with $r_{i}=l_{i+1}$ for adjacent clusters. We then arrange for the denoising tasks to become progressively harder as we move from cluster $N$ to cluster $1$, ensuring $l_{i}<l_{i+1}$, with $l_{1}=0$ and $r_{N}=T$. During training, we introduce these clusters sequentially, starting from the least challenging (the $N$-th cluster) and advancing towards the most challenging (the $1$-st cluster), through $N$ distinct phases. Specifically, in the $n$-th stage, we jointly train the model on clusters $\{C_{N},C_{N-1},\ldots,C_{N-(n-1)}\}$. After this $N$-phase of curriculum learning, the model continues to learn across the entire range of timesteps, progressively converging towards a point of stability.
@@
The next consideration involves determining the boundaries for each cluster, $l_{i}$ and $r_{i}$. A straightforward approach is to uniformly divide the entire timestep interval $[0,T]$ as $C_{i}=[\frac{(i-1)\cdot T}{N},\frac{i\cdot T}{N}]$ for $i=1,2,\ldots,N$. However, this method does not account for variations in noise levels across different timesteps. Therefore, to address this issue more effectively, we adopt an SNR-based interval clustering technique as used in [11], which aligns the clustering with the actual changes in noise levels, potentially enhancing curriculum learning adaptability to varying noise conditions.
@@
In contrast, for EDM [23] which operates based on the noise level $\sigma$ rather than the timestep $t$, and where $\sigma$ is sampled from a log-normal distribution such that $\log(\sigma)\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$ during training, our clustering strategy for timesteps cannot be directly transposed. Given the log-normal distribution of $\sigma$, dividing it directly is impractical because $\sigma$ can extend over a wide range of values. Instead, we adapt our clustering approach to suit the log-normal characteristics by defining noise level clusters $C_{i}$. Specifically, we delineate $C_{i}=[\mathrm{\Phi}^{-1}(\frac{i-1}{N}),\mathrm{\Phi}^{-1}(\frac{i}{N})]$, where $\mathrm{\Phi}^{-1}$ is the inverse cumulative distribution function (quantile function) of the Gaussian distribution $\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$. This method segments the noise levels into intervals by reflecting their probabilistic distribution.
Algorithm 1 Curriculum Learning Input: Curriculum $\{C_{i}\}_{i=1}^{N}$, Pacing function $g$, Maximum patience $\tau_{\text{max}}$, Loss function $f$, Curriculum index $I_{\text{cur}}=N$, Best loss $L_{\text{best}}=\infty$, Model $M_{\theta}$ while $I_{\text{cur}}>0$ do     # Mini-batch sampling     $X\sim C_{I_{\text{cur}}}$     # Calculate Loss     $L_{\text{cur}}=f(M_{\theta}(X))$     # Update model     $\theta$ = $\theta$ $\nabla_{\theta}{L_{\text{cur}}}$     # Pacing function     $\tau_{\text{cur}},I_{\text{next}}=g(L_{\text{cur}},L_{\text{best}},\tau_{\text{cur}},\tau_{\text{max}},I_{\text{cur}})$     # Update curriculum     if $I_{\text{cur}}\neq I_{\text{next}}$ then        $I_{\text{cur}}=I_{\text{next}}$        $L_{\text{best}}=\infty$    # Update best loss     else if $L_{\text{cur}}<L_{\text{best}}$ then        $L_{\text{best}}=L_{\text{cur}}$     end if end while
Algorithm 2 Pacing Function Input: Current loss $L_{\text{cur}}$, Best loss $L_{\text{best}}$, Current patience $\tau_{\text{cur}}$, Maximum patience $\tau_{\text{max}}$, Current curriculum index $I_{\text{cur}}$ Output: Updated patience, Updated curriculum index # Reset patience if $L_{\text{cur}}<L_{\text{best}}$ then     return 0, $I_{\text{cur}}$ else     # Proceed to next curriculum     if $\tau_{\text{cur}}+1>\tau_{\text{max}}$ then        return 0, $I_{\text{cur}}-1$    # Increase patience     else        return $p_{\text{cur}}+1,I_{\text{cur}}$     end if end if
@@ 5.2 Pacing Strategy of Curriculum
To effectively train the diffusion model according to the provided curriculum design, it is crucial to define a suitable pacing function for determining the transition of each $N$ distinct curriculum stage. Training for a fixed number of iterations for each curriculum is the simplest implementation (We also contain this method in experiments as ‘NaiveCL’ in Section 6.2). However, the convergence rate of each curriculum varies significantly, as demonstrated in Fig. 1. Hence, we propose adopting an adaptive number of iterations for each curriculum, akin to the varied exponential pacing approach explored by Hacohen et al. [13]. The proposed pacing function considers the loss of a mini-batch and maximum patience. If the loss does not improve within the maximum patience iterations, the current curriculum is terminated, and the subsequent curriculum is initiated. Here, the maximum patience is a fixed hyper-parameter and the detailed process is outlined in Algorithm 2. The overall curriculum learning procedure is presented in Algorithm 1.
@@ 6.1 Experimental Setup
Here, we provide experimental setups concisely due to space constraints. Detailed setups are presented in Appendix C.
6.1.1 Evaluation protocols.
For our comprehensive evaluation of various methods, we employed three distinct image-generation tasks: 1) Unconditional generation: We utilized the FFHQ dataset [23], which comprises 70,000 training images of human faces. 2) Class-conditional generation: For this task, we employed two datasets: CIFAR-10 [28], which includes 60,000 training images, and ImageNet [7], containing 1,281,167 training images. In this setup, we applied classifier-free guidance [19]. 3) Text-to-Image generation: We used the MS-COCO dataset [33] , which contains 82,783 training images and 40,504 validation images. Each image in this dataset is accompanied by five text captions, facilitating the text-to-image generation task.
6.1.2 Evaluation metrics.
To evaluate the performance of models, we utilized three metrics: FID [16], IS [49], and Precision/Recall [30]. Specifically, we applied FID and IS to measure sample quality, while Precision is used to assess quality further and Recall was utilized to evaluate the diversity of the generated samples in ImageNet setup. In other datasets, we employed FID to evaluate sample quality.
6.1.3 Target models.
We employed two exemplary diffusion architectures for experiments: DiT [42], which integrates latent diffusion models [46] with Transformer architectures [58], and EDM [22], which focuses on pixel-level diffusion utilizing UNet-based architectures [47]. For the text-to-image generation, we incorporated a CLIP text encoder [44] as described in DTR [41].
@@ 6.2 Comparative Results
In this section, we assess the effectiveness of our curriculum-based training approach. For a thorough comparison, we examine three distinct training variants, with further details provided in Appendix C: To achieve this, we compare three variants of training: 1) Vanilla: This term refers to diffusion models trained using conventional methods without any curriculum learning strategies; 2) NaiveCL: In this variant, we incorporate a basic curriculum learning strategy, which simply repeats the same number of iterations for each stage across an $N$-stage process and does not employ SNR-based clustering; 3) Ours: This denotes our proposed curriculum approach, which is designed to enhance the training process of diffusion models by systematically structuring the learning stages.
6.2.1 Quantitative evaluation.
As shown in Table 1, our approach was validated on various architectures—DiT [42] and EDM [22]— and tasks including unconditional, class-conditional, and text-to-image generation. To further contextualize our method, we evaluated the performance of NaiveCL, which does not leverage the SNR-based clustering and pacing function proposed in our approach. Our observations indicate that NaiveCL fails to consistently achieve improved performance, as it lacks the capability to incorporate task difficulty. In contrast, our method explicitly integrates the difficulty of denoising tasks, as detailed in Section 4. Consequently, our approach consistently demonstrates significant performance enhancements across all metrics on four datasets: FFHQ [23], ImageNet [7], CIFAR-10 [28], and MS-COCO [33], demonstrating that our approach works well data- and model-agnostic.
Latent-Level Diffusion Model.
@@
Table 1: We evaluated unconditional image generation on FFHQ with DiT-B and EDM, class-conditional image generation on ImageNet and CIFAR10 with DiT-L and EDM, respectively, and text-conditional image generation on MS-COCO with DiT-B. Note that our curriculum learning for diffusion models improves substantial performance without any additional parameters.
6.2.2 Qualitative evaluation.
Due to space constraints, we illustrate a detailed collection of generated examples in Appendix D. In summary, our curriculum methodology demonstrates a notable enhancement in the quality of the images produced, especially when compared to naive training.
@@ 6.3 Analysis
To elucidate the effectiveness of our curriculum approach, we hereby present a series of analytical studies. All the analysis is conducted by using the DiT-B model on the ImageNet dataset.
6.3.1 Effects of $N$ and $\tau$.
@@
We examined the robustness of the proposed curriculum training with respect to hyper-parameters: the number of clusters $N$ and the maximum patience $\tau$. As shown in Fig. 3  our method consistently outperforms the vanilla model, and the best result is observed at $N=20,\tau=200$. It shows that as $\tau$ increases, it may lead to overtraining due to excessive iterations for each task, whereas as $\tau$ decreases, curriculums may not be sufficiently trained. Furthermore, when the entire range of timesteps is finely partitioned (i.e., with an increase in $N$), each cluster becomes excessively granular, resulting in suboptimal performance. Conversely, with a decrease in $N$, tasks that should be in distinct clusters are learned together, forming a coarser cluster, which also leads to suboptimal outcomes. We utilized the optimal hyper-parameters for all subsequent experiments.
6.3.2 Effects of Curriculum Design.
@@
Table 2: Comparative results on various curriculum designs.
In our curriculum design, we initially partitioned the entire set of timesteps into $N$ clusters using SNR-based clustering, organizing the curriculum from easy to hard clusters. To thoroughly assess the impact of each component, we conducted the ablation study as shown in Table. 2. Firstly, we investigated the effect of curriculum learning via comparison with an anti-curriculum approach [13], which progresses from hard to easy clusters. While both training methods that consider task order (b2, c2) appear to enhance performance compared to vanilla training (a), anti-curriculum training cannot consistently guarantee performance improvement. To specifically assess the impact of anti-curriculum learning, we evaluate the performance of anti-curriculum learning without SNR-based clustering (b1), resulting in a performance decrease compared to vanilla. This can be attributed to the deviation from the observed appropriate task sequence, with the improvement seen in (b2) attributed to SNR-based clustering. In contrast, the proposed curriculum learning method (c1, c2) consistently exhibited performance improvement even with the uniformly partitioned clusters. Besides, with the finding that utilizing SNR-clustering was more effective, the clustering with the actual changes in noise levels enhanced the curriculum learning adaptability. We believe that by embracing more advanced clustering methods, we could gain further performance improvements.
6.3.3 Visualization of curriculum.
@@
To gain deeper insights into the functioning of our curriculum pacing, we plotted the loss metrics against the curriculum phases, as illustrated in Fig. 5. During the curriculum training, tasks progressively transition from the easiest to the most challenging, with varying amounts of iterations for each task based on the pacing function. The training loss decreased during each curriculum phase but increased after curriculum changes via the pacing function due to the inclusion of a newly added task in the updated curriculum. In addition, as $\tau$ increases, it is confirmed that the curriculum phase slowly changes.
6.3.4 Analysis on convergence speed.
As demonstrated in previous works [3, 13], the adoption of curriculum learning can lead to faster convergence in model performance. To illustrate the efficacy of our approach in this regard, we plotted the FID, IS, precision, and recall calculated over 10,000 samples across the training iterations, as depicted in Fig. 5. We observed the models trained through the proposed curriculum learning method converge faster than vanilla models, regardless of evaluation metrics. Note that our approach does not need any additional parameters and training iterations.
6.3.5 Robustness on various sizes of models
To verify the generalizability of our method across different model sizes, we evaluated the performance gains achieved using our curriculum learning approach on various scales of the DiT model: DiT-S (small), DiT-B (base), and DiT-L (large). Table. 4 shows that the proposed curriculum learning for diffusion model improves the performance regardless of the model size. Moreover, it is notable that larger models exhibit a more substantial performance enhancement: DiT-S has improved by 8% in terms of FID, while DiT-B and DiT-L have shown improvements of 24% and 27%, respectively. These findings validate the efficacy of our curriculum approach across a diverse range of model sizes, underscoring its robust generalizability.
6.3.6 Additional experimental results
Due to limited space, we present additional experimental results in Appendix E. These results also support the effectiveness of our method, emphasizing the importance of curriculum approaches in diffusion training.
6.4 Orthogonality of Our Curriculum Approach
@@
Table 4: Note that the curriculum learning is compatible with the previous works such as loss weighting (MinSNR) and architecture (DTR) study which, specified the multi-task learning for diffusion model.

### v2

@@ 5.1 Design of Curriculum Learning in Diffusion Models
Figure 3: The overview of our curriculum learning approach for diffusion models. (Left) We divide the timesteps into $N$ clusters, ${C_{1},...,C_{N}}$, with the difficulty of denoising tasks increasing from $C_{N}$ (easiest) to $C_{1}$ (hardest). (Right) As the curriculum progresses, learning accumulates harder task clusters, gradually increasing task difficulties.
@@
As we observed in Section 4, difficulties in denoising tasks increase as $t$ gets smaller. To utilize an easy-to-hard curriculum learning approach, we first divide the entire range of timesteps into $N$ clusters, denoted as $\{C_{i}\}_{i=1}^{N}$, where each cluster $C_{i}$ spans an interval $[l_{i},l_{i+1}]$, ensuring $l_{i}<l_{i+1}$, with $l_{1}=0$ and $l_{N+1}=T$, as shown on the left side of Fig. 3. The curriculum for training is constructed by regarding these task clusters as unit tasks, starting from the least challenging (the $N$-th cluster $C_{N}$) and advancing towards the most difficult (the first cluster $C_{1}$), through $N$ distinct stages. Specifically, in the $n$-th curriculum stage, we jointly train the model with denoising tasks sampled from the clusters $\bigcup_{j=N-(n-1)}^{N}C_{j}$ as illustrated in the right side of Fig. 3. The transition of the curriculum stages is determined by the pacing function, which will be discussed in the next section. After completing these $N$-stages of curriculum learning, the model continues to learn across the entire range of timesteps, $\bigcup_{j=1}^{N}C_{j}$, same as standard diffusion training.
@@
The next consideration involves determining the boundaries for each cluster $l_{i}$. A straightforward approach is to uniformly divide the entire timestep interval $[0,T]$ as $C_{i}=[\frac{(i-1)\cdot T}{N},\frac{i\cdot T}{N}]$ for $i=1,2,\ldots,N$. However, this method does not account for variations in noise levels across different timesteps. Therefore, to address this issue more effectively, we adopt an SNR-based interval clustering technique as used in [11], which aligns the clustering with the actual changes in noise levels, potentially enhancing curriculum learning adaptability to varying noise conditions.
@@
For EDM [24] which operates based on the noise level $\sigma$ rather than the timestep $t$, and where $\sigma$ is sampled from a log-normal distribution such that $\log(\sigma)\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$ during training, our clustering strategy for timesteps cannot be directly transposed. Given the log-normal distribution of $\sigma$, dividing it directly is impractical because $\sigma$ can extend over a wide range of values. Instead, we adapt our clustering approach to suit the log-normal characteristics by defining noise level clusters $C_{i}$. Specifically, we delineate $C_{i}=[\mathrm{\Phi}^{-1}(\frac{i-1}{N}),\mathrm{\Phi}^{-1}(\frac{i}{N})]$, where $\mathrm{\Phi}^{-1}$ is the inverse cumulative distribution function (quantile function) of the Gaussian distribution $\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$. This method segments the noise levels into intervals by reflecting their probabilistic distribution.
@@ 5.2 Pacing Strategy of Curriculum
To effectively train the diffusion model according to the provided curriculum design, it is crucial to define a suitable pacing function for determining the transition of each $N$ distinct curriculum. Training for a fixed number of iterations for each curriculum stage is the simplest implementation (We also contain this method in experiments as ‘NaiveCL’ in Section 6.2). However, the convergence rate of each curriculum phase varies significantly, as demonstrated in Fig. 1. Hence, we propose adopting an adaptive number of iterations for each curriculum, akin to the varied exponential pacing approach explored by Hacohen et al. [13]. Our pacing function utilizes the training loss to determine transition moments and transitions to the next stage occur when the training loss converges at the current stage. Specifically, we introduce the maximum patience iteration $\tau$, and if the loss does not improve consecutively for $\tau$, the current curriculum stage is terminated, and the subsequent curriculum stage is initiated. Here, the maximum patience is a fixed hyper-parameter, and the detailed process and overall curriculum learning procedure are outlined in Algorithm 1 and 2 in Appendix D, respectively.
@@ 6.1 Experimental Setup
Here, we provide experimental setups concisely. Detailed setups are presented in Appendix E.
Evaluation protocols.
For our comprehensive evaluation of various methods, we employed three distinct image-generation tasks: 1) Unconditional generation with the FFHQ dataset [24], 2) Class-conditional generation with CIFAR-10 [28] and ImageNet [7] datasets, and 3) Text-to-Image generation with MS-COCO dataset [33]. In 2) and 3) setups, we applied classifier-free guidance [19].
Target models.
We employed three exemplary diffusion architectures for experiments: DiT [44], which integrates latent diffusion models [48] with Transformer architectures [60] parameterized as $\epsilon$-prediction, EDM [22], which focuses on pixel-level diffusion utilizing UNet-based architectures [49] parameterized as $F$-prediction, and SiT [38] for score- and velocity-prediction. For the text-to-image generation, we incorporated a CLIP text encoder [46] as described in DTR [43].
@@ 6.2 Comparative Results
In this section, we assess the effectiveness of our curriculum-based training approach. For a thorough comparison, we examine three distinct training variants, with further details provided in Appendix E: To achieve this, we compare three variants of training: 1) Vanilla: This term refers to diffusion models trained using conventional methods without any curriculum learning strategies; 2) NaiveCL: In this variant, we incorporate a basic curriculum learning strategy, which simply repeats the same number of iterations for each stage across an $N$-stage process and does not employ SNR-based clustering; 3) Ours: This denotes our proposed curriculum approach, which is designed to enhance the training process of diffusion models by systematically structuring the learning stages.
Quantitative evaluation.
We quantitatively validate the effectiveness of our methods with various architectures-DiT [44], EDM [22], and SiT [38]- and tasks including unconditional, class-conditional, and text-to-image generation. Table 1 shows the results, confirming two empirical observations: 1) NaiveCL fails to consistently achieve improved performance compared to Vanilla, and 2) our approach outperforms both NaiveCL and Vanilla. Regarding the first observation, NaiveCL shows inconsistent improvements due to its lack of robust adaptation on incorporating task difficulties in various task conditions. In contrast, our method demonstrates superior performance across all scenarios by improving the clustering and pacing of curriculums. Consequently, our approach consistently achieves significant performance enhancements across all metrics on four datasets: FFHQ [24], ImageNet [7], CIFAR-10 [28], and MS-COCO [33], illustrating its effectiveness regardless of data or model used.
Table 1: We evaluated unconditional image generation on FFHQ with DiT-B, EDM, and SiT-B, class-conditional image generation on ImageNet and CIFAR10 with DiT-L and EDM, respectively, and text-conditional image generation on MS-COCO with DiT-B. Note that our curriculum learning for diffusion models improves substantial performance without any additional parameters.
$\epsilon$-prediction
@@
Velocity-prediction
Model FFHQ 256$\times$256
[…]
SiT (Vanilla) 7.44
SiT + NaiveCL 7.69
SiT + Ours 6.95
Score-prediction
Model FFHQ 256$\times$256
[…]
SiT (Vanilla) 9.64
SiT + NaiveCL 9.77
SiT + Ours 9.15
Qualitative evaluation.
Due to space constraints, we illustrate a detailed collection of generated examples in Appendix F. In summary, our curriculum methodology demonstrates a notable enhancement in the quality of the images produced, when compared to NaiveCL and Vanilla.
@@ 6.3 Analysis
To elucidate our curriculum approach’s effectiveness, we present a series of analytical studies. All the analysis is conducted by using the DiT-B model on the ImageNet dataset.
@@
Effects of $N$ and $\tau$.
We examined the robustness of the proposed curriculum training with respect to hyper-parameters: the number of clusters $N$ and the maximum patience $\tau$. As shown in Fig. 4, our method consistently outperforms the vanilla model, and the best result is observed at $N=20,\tau=200$. It shows that as $\tau$ increases, it may lead to overtraining due to excessive iterations for each task, whereas as $\tau$ decreases, curriculums may not be sufficiently trained. Furthermore, when the entire range of timesteps is finely partitioned (i.e., with an increase in $N$), each cluster becomes excessively granular, resulting in suboptimal performance. Conversely, with a decrease in $N$, tasks that should be in distinct clusters are learned together, forming a coarser cluster, which also leads to suboptimal outcomes. Overall, our method outperforms vanilla training across a range of hyperparameters, demonstrating the robustness of our approach.
Effects of Curriculum Design.
Table 2: Comparative results on various curriculum designs.
@@
In our curriculum design, we initially partitioned the entire set of timesteps into $N$ clusters using SNR-based clustering, organizing the curriculum from easy to hard clusters. To thoroughly assess the impact of each component, we conducted the ablation study as shown in Table 2. Firstly, we investigated the effect of curriculum learning via comparison with an anti-curriculum approach [13], which progresses from hard to easy clusters, unlike conventional curriculum learning. While both training methods in (b2) appear to enhance performance compared to vanilla training (a), anti-curriculum training cannot consistently guarantee performance improvement concerning the curriculum design as shown in (b1). In contrast, the proposed curriculum learning method (c1, c2) consistently exhibited performance improvement even with the uniformly partitioned clusters. Besides, with findings that utilizing SNR-clustering was more effective, clustering with the actual changes in noise levels enhanced the curriculum learning adaptability.
Visualization of curriculum.
To gain deeper insights into the functioning of our curriculum pacing, we plotted loss metrics against curriculum phases, as illustrated in Fig. 6. During the curriculum training, tasks progressively transition from the easiest to the most challenging, with varying amounts of iterations for each task based on the pacing function. The training loss decreased during each curriculum phase but increased after curriculum changes via the pacing function due to the inclusion of a newly added task in the updated curriculum. Additionally, as $\tau$ increases, the curriculum phases change more gradually, highlighting the role of $\tau$ in controlling the pace of curriculum transitions.
Analysis on convergence speed.
As demonstrated in previous works [3, 13], the adoption of curriculum learning can lead to faster convergence in model performance. To illustrate the efficacy of our approach in this regard, we plotted the FID, IS, precision, and recall calculated over 10,000 samples across the training iterations, as depicted in Fig. 6. We observed the models trained through the proposed curriculum learning method converge faster than vanilla models, regardless of evaluation metrics. Notably, our approach achieves these improvements without requiring additional parameters or training iterations, thereby significantly saving time and computational resources.
@@
Effectiveness on various sizes of models
To verify the generalizability of our method across different model sizes, we evaluated the performance gains achieved using our curriculum learning approach on various scales of the DiT model: DiT-S (small), DiT-B (base), and DiT-L (large). Table. 4 shows that the proposed curriculum learning for diffusion model improves the performance regardless of the model size. Moreover, it is notable that larger models exhibit a more substantial performance enhancement: DiT-S improved by 8% in terms of FID, while DiT-B and DiT-L showed improvements of 24% and 27%, respectively. These findings validate the efficacy of our curriculum approach across a diverse range of model sizes, underscoring its generalizability to various model parameters.
Orthogonality of Our Curriculum Approach
@@
Additional experimental results
Due to limited space, we present additional experimental results in Appendix G. These results also support the effectiveness of our method, emphasizing the importance of curriculum approaches in diffusion training.
@@
DiT-S/2 43.30 33.63 0.42 0.54
DiT-S/2 + Ours 39.66 36.57 0.44 0.54
DiT-B/2 30.27 60.06 0.55 0.52
DiT-B/2 + Ours 22.96 75.98 0.62 0.52
DiT-L/2 11.18 146.95 0.75 0.47
DiT-L/2 + Ours 8.18 186.37 0.79 0.47
Table 4: Note that the curriculum learning is compatible with the previous works such as the loss weighting (MinSNR) and architecture (DTR) study which, specified the multi-task learning for diffusion model.
Class-Conditional ImageNet 256$\times$256.
DiT-B/2 DiT-B/2 + Ours
@@
Vanilla 30.27 60.06 0.55 0.52 22.96 75.98 0.62 0.52
MinSNR [14] 21.88 88.12 0.63 0.49 19.36 101.35 0.67 0.49
DTR [43] 15.77 89.89 0.68 0.52 15.33 91.39 0.68 0.52